Could someone please show me step by step how to Write a quadratic equation in the variable x having the given numbers as solutions. Type the eqyation in standard form, ax^2+bx+c=0. solution -√5,6√5
Answer by stanbon(75887)

Write a quadratic equation in the variable x having the given numbers as solutions. Type the eqyation in standard form, ax^2+bx+c=0. solution -√5,6√5
----
Note: If "a" is a solution, x-a is a factor.
----
Your problem:
y = (x+sqrt(5))(x-6sqrt(5))
---
y = x^2 +(sqrt(5)-6sqrt(5))x -6*5
----
y = x^2 + (sqrt(5)-6sqrt(5))x -30
=====================================
Cheers,
Stan H.
===========